# Tony Lindeberg

### v5

@@ 3.3 Orientation selectivity histograms for the generalized idealized models of complex cells
Based on these results, we therefore propose to (i) include an explicit variability over the degree of elongation of the receptive fields, (ii) include third- and fourth-order models of simple cells in addition to previous use of first- and second-order models of simple cells and (iii) integrate non-linear transformations of such receptive field responses over extended regions in image space, when modelling the functional properties of complex cells.
4 Summary of discussion
@@
We have presented a set of three new integrated affine quasi quadrature measures to model the functional properties of complex cells, and analyzed their orientation selectivity properties, based on the assumption that the receptive field shapes span a variability over the degree of elongation of the simple cells that form the input to these models of complex cells.
@@
Specifically, by qualitative comparisons with the biological results by Nauhaus et al. (2008), regarding a significant variability in orientation selectivity properties of biological neurons from wide to narrow orientation selectivity properties, and to Goris et al. (2015), regarding orientation selectivity histograms over the resultant $|R|$, we have found that these results are consistent with a previously proposed hypothesis in Lindeberg (2023) further investigated in Lindeberg (2024a) that the receptive fields in the primary visual cortex should span a significant variability in the degree of elongation of the receptive fields. In this respect, the results are consistent with what becomes a natural consequence of stating desirable properties of an idealized vision system, that the receptive fields should be covariant under the natural geometric image transformations. In such a context, covariance properties of the family of receptive fields enable more accurate estimates of cues to the 3-D structure of the world, as the image data used as input to the vision system undergo significant variabilities, as caused by variations of the viewing conditions, such as the distance and the viewing direction between objects in the world and the observer.
Specifically, by simulating the orientation selectivity histograms that result from the presented new integrated affine quasi quadrature measures ${\cal Q}_{\varphi,12,\text{int}}L$, ${\cal Q}_{\varphi,1234,\text{int}}L$ and ${\cal Q}_{\varphi,34,\text{int}}L$ according to (11), (12) and (13), when combined with an assumption of a uniform distribution over the logarithm of the scale parameter ratio $\kappa=\sigma_{2}/\sigma_{1}$ of the receptive fields, we have found that the extensions of the previous pointwise quasi quadrature measure ${\cal Q}_{\varphi,12,\text{int}}L$ according to (3) offer ways of changing the shapes of the predicted orientation selectivity histograms regarding both how uniform the predicted histograms will be in relation to the previously recorded biological orientation selectivity histograms by Goris et al. (2015) and regarding the span of values of the resultant $|R|$ they cover.
@@
Let us finally remark that it should most likely be the case that the non-linear behaviour of complex cells may be more complex than the computational mechanisms used in the proposed idealized models in terms of integrated affine quasi quadrature measures. The overall purpose with this work is on the other hand to demonstrate that the orientation selectivity properties of the proposed idealized models of complex cells can be analyzed with a structurally similar methodology as used for probing the orientation selectivity properties of biological receptive fields. From this viewpoint, the comparison to the biological orientation selectivity histogram accumulated by Goris et al. (2015) is specifically to show that the gross behaviour in terms of an underlying distribution of receptive field shapes of different elongation can be used to reflect gross properties of the biological measurements. Our intention in this respect is to stimulate further tests of more complex computational models of complex cells, based on assuming distributions of the underlying receptive field shapes over the degree of elongation.
From the viewpoint of quasi quadrature models to be used for addressing computational tasks in computer vision, it should on the other hand also be mentioned that a hierarchical network constructed by applying a substantially simplified quasi quadrature model of complex cells in cascade can lead to quite reasonable results on computer vision benchmarks (Lindeberg 2020); see also the closely related work by Riesenhuber and Poggio (1999), Serre et al. (2007) and Pant et al. (2024), who construct hierarchical networks for computer vision tasks from sets of idealized models of simple and complex cells coupled in cascade. These studies in computational vision do thus demonstrate that computational mechanisms structurally closely related to the affine quasi quadrature measures proposed and studied in this work can support spatial recognition tasks on visual data.

### v7

@@ Results
Based on these results, we therefore propose to (i) include an explicit variability over the degree of elongation of the receptive fields, (ii) include third- and fourth-order models of simple cells in addition to previous use of first- and second-order models of simple cells and (iii) integrate non-linear transformations of such receptive field responses over extended regions in image space, when modelling the functional properties of complex cells. Specifically, we propose that the items (i) and (iii) above should be of much wider generality than restricted to computational models based on affine quasi quadrature measures, and thereby also apply to other families of computational models of complex cells.
Explicit predictions for further modelling of complex cells
More explicitly, based on the above results, we can thus state the following general predictions:
Prediction 1 (Flexibility in elongation would lead to better approximation properties when modelling individual complex cells): Given a sufficiently large set of biological complex cells, if the receptive fields of those are modelled by computational models, then such models that involve a flexibility in the degree of elongation of the underlying computational primitives would lead to better approximation properties than for computational models that do not involve a flexibility in the degree of elongation.
Prediction 2 (Variability in elongation over populations of complex cells): If the receptive fields sufficiently large set of complex cells are modelled by computational models that involve a flexibility in the degree of elongation of the underlying computational primitives, then such model fitting would lead to a substantial variability in the degree of elongation of the receptive fields over a sufficiently large population of modelled biological complex cells.
Prediction 3 (Spatial integration as a computational mechanism in complex cells): Given a sufficiently large set of biological complex cells, if their receptive fields are modelled by computational models, then models that are based on spatial integration of underlying computational primitives would lead to better approximation properties than computational models that do not comprise any spatial integration.
Specifically, we propose that these predictions could be explored and influence further modelling of complex cells in terms of mathematically based image primitives.
Explicit suggestion to extension of methodology for experimentally probing the orientation selectivity of complex cells that may comprise a variability in the degree of elongation between different visual neurons.
Furthermore, to allow for better distinctions between the validity of different types of computational models for complex cells, we propose to extend the experimental methodology for measuring the orientation selectivity properties of biological neurons to instead of (i) choosing a single angular frequency for the probing sine wave for generating the visual stimuli for each spatial image orientation, instead (ii) performing a simultaneous two-parameter variation over both the angular frequency and the image orientation. In such a way, the experimental data ought to be better suited for handling biological receptive fields with a variability in elongation. The reason for this, is that the generated visual stimuli would then better probe the dependency of two characteristic inherent spatial scales of the biological receptive fields compared to using a single inherent spatial scale for the visual stimuli.
As described in more detail in Lindeberg [16] Section 7, the choice of the angular frequency for sine wave for probing the orientation selectivity properties of visual neurons can significantly affect the shapes of the resulting orientation selectivity curves, and should thus warrant specific consideration.
Summary and discussion
@@
We have presented a set of three new integrated affine quasi quadrature measures to model the functional properties of complex cells, and analyzed their orientation selectivity properties, based on the assumption that the receptive field shapes should span a variability over the degree of elongation of the simple cells that form the input to these models of complex cells.
@@
Specifically, by qualitative comparisons with the biological results by Nauhaus et al. [22], regarding a significant variability in orientation selectivity properties of biological neurons from wide to narrow orientation selectivity properties, and to Goris et al. [15], regarding orientation selectivity histograms over the resultant $|R|$, we have found that these results are consistent with a previously proposed hypothesis in Lindeberg [10] further investigated in Lindeberg [14] that the receptive fields in the primary visual cortex should span a significant variability in the degree of elongation of the receptive fields. In this respect, the results are consistent with what becomes a natural consequence of stating desirable properties of an idealized vision system, that the receptive fields should be covariant under the natural geometric image transformations. In such a context, covariance properties of the family of receptive fields enable more accurate estimates of cues to the 3-D structure of the world, as the image data used as input to the vision system undergo significant variabilities, as caused by variations of the viewing conditions, such as the distance and the viewing direction between objects in the world and the observer.
Specifically, by simulating the orientation selectivity histograms that result from the presented new integrated affine quasi quadrature measures ${\cal Q}_{\varphi,12,\text{int}}L$, ${\cal Q}_{\varphi,1234,\text{int}}L$ and ${\cal Q}_{\varphi,34,\text{int}}L$ according to (11), (12) and (13), when combined with an assumption of a uniform distribution over the logarithm of the scale parameter ratio $\kappa=\sigma_{2}/\sigma_{1}$ of the receptive fields, we have found that the extensions of the previous pointwise quasi quadrature measure ${\cal Q}_{\varphi,12,\text{int}}L$ according to (3) offer ways of changing the shapes of the predicted orientation selectivity histograms regarding both how uniform the predicted histograms will be in relation to the previously recorded biological orientation selectivity histograms by Goris et al. [15] and regarding the span of values of the resultant $|R|$ they cover.
@@
Let us finally remark that it should most likely be the case that the non-linear behaviour of complex cells may be more complex than the computational mechanisms used in the proposed idealized models in terms of integrated affine quasi quadrature measures. The overall purpose with this work is on the other hand to demonstrate that the orientation selectivity properties of the proposed idealized models of complex cells can be analyzed with a structurally similar methodology as used for probing the orientation selectivity properties of biological receptive fields. From this viewpoint, the comparison to the biological orientation selectivity histogram accumulated by Goris et al. [15] is specifically to show that the gross behaviour in terms of an underlying distribution of receptive field shapes of different elongation can be used to reflect gross properties of the biological measurements. Our intention in this respect is to stimulate further tests of more complex computational models of complex cells, based on assuming distributions of the underlying receptive field shapes over the degree of elongation. The set of more explicit predictions in the section “Explicit predictions for further modelling of complex cells” are specifically aimed at providing a guide to such further research.
The proposed extension of the methodology for characterizing the orientation selectivity of visual neurons in the section “Explicit suggestion to extension of methodology for experimentally probing the orientation selectivity of complex cells that may comprise a variability in the degree of elongation between different visual neurons.”, by instead of performing a one-parameter variation over the orientation of the probing sine wave instead performing a two-parameter variation over both the angular frequency and the orientation of the probing sine wave, is intended to provide richer experimental data that could better distinguish between the explanatory properties of different types of computational models of complex cells.
From the viewpoint of quasi quadrature models to be used for addressing computational tasks in computer vision, it should on the other hand also be mentioned that a hierarchical network constructed by applying a substantially simplified quasi quadrature model of complex cells in cascade can lead to quite reasonable results on computer vision benchmarks (Lindeberg [89]); see also the closely related work by Riesenhuber and Poggio [102], Serre et al. [103] and Pant et al. [104], who construct hierarchical networks for computer vision tasks from sets of idealized models of simple and complex cells coupled in cascade. These studies in computational vision do thus demonstrate that computational mechanisms structurally closely related to the affine quasi quadrature measures proposed and studied in this work can support spatial recognition tasks on visual data.